OpenStudy (anonymous):

A policeman on a motorcycle is pursuing a car that is speeding at 115 kmper hour. The policeman is 6 km behind the car and is traveling 130 kmper hour. How long will it be before the policeman overtakes the car?

OpenStudy (anonymous):

distance = speed * time difference in speed = 130- 115 = 15 Km/hr distance between them = 6 Km time taken before the police overtakes = distance bewteen them/diferrence in speed = 6 / 15 = 0.4 hrs or 24 min